What mass of AlF3 is produced when 168 grams of Al is consumed?


1
Expert's answer
2021-06-11T01:58:34-0400

Solution

Al(s)+F2(g)➡️AlF3(s)


1mole Al produces 1 mole AlF3


RAM of Al=27g, F=19g

RFM of AlF3=84g

No of moles = "\frac{Mass}{RFM}=\frac{168}{27}"

= 6.2 moles

6.2 moles of Al=6.2 moles of AlF3

Mass= Number of moles"\times RFM"

"6.2\times84=520.8"


Mass of AlF3 produces is 520.8g